# Swarm Intelligence Enhanced Reasoning: A Density-Driven Framework for LLM-Based Multi-Agent Optimization

### v1

@@ 2.1 Multi-Agent Systems
While the MAD scheme has shown effectiveness, previous systematic evaluations zhang2025if revealed that many MAD methods perform even worse than CoT methods. The advantage of solution diversity introduced by heterogeneous agent roles does not always hold, as the agents ultimately rely on the same underlying large language model, which leads to high similarities of the solutions generated by agents. Furthermore, the similarity between the solutions gradually increases during the collaboration process, since the LLM is progressively influenced by the accumulated context from historical interaction information. This factor makes it difficult to find the optimal results in the final despite the use of a collaborative scheme. Different from them, we propose integrating swarm intelligence algorithms into the reasoning process to enhance the discovery of optimal solutions. We conceptualize the LLM’s reasoning process as an exploration of the solution space and adopt heuristic algorithms to efficiently identify optimal solutions.
@@ 4 Methodology
Swarm intelligence algorithms typically explore the solution space via individual exploration and collective cooperation to identify the global optimum. This objective is analogous to that of LLM-based agents when undertaking reasoning tasks. Consequently, we proposed a novel paradigm named ASI, which conceptualized the LLM’s reasoning process as an individual’s search for the global optimum within a solution space. Specifically, for a given problem $Q$, it owns a solution space $S$. Each dimension corresponds to a token, with its value derived from the LLM’s vocabulary. Every point within this solution space denotes a solution in the form of a reasoning path, and the essence of LLM reasoning is to generate such a path. Therefore, we applied swarm intelligence algorithms to LLM reasoning, which leverages the evaluator $E$ to evaluate the quality of the reasoning paths and guide the search, iteratively refining reasoning paths to efficiently and effectively uncover the global optima. Furthermore, based on the ASI paradigm, we propose the Swarm Intelligence Enhancing Reasoning (SIER) framework, which enhances solution space exploration by expanding the diversity of the search paths. This section first presents an overview of the SIER framework and detailed explanations of its key processes.
@@ B.1 Analysis of MAD Framework
Although the existing MAD framework is based on the hypothesis that LLMs can play different roles under various human prompts, unlike simple CoT, heterogeneous prompts can help agents to collectively diverge their thinking and decide the final answer, which attempts to achieve two goals:
Brainstorming: Broaden the width of exploration of the solution space, and obtain a variety of candidate solutions when facing complex problems.
@@
However, existing evaluation work suggests that MAD frameworks do not perform as well as expected with respect to the above two goals. [45] experimentally found that many MAD frameworks perform worse than single-point methods and that self-consistency (i.e., by sampling the CoT results multiple times and voting for the best result) outperforms MAD frameworks significantly, as shown in Table 3.
@@
Besides, to analyze the reasons for poor MAD performance, [36] identified two key problems that are relevant to the goal Selection: (1) Judgment Error: This occurs when the judger decides the final answer. If the answers vary between agents, the judger may choose the wrong option as the final verdict. This is especially true when the decision is made in a tie. (2) Wrong Answer Propagation: This error occurs when an agent, influenced by input from others, deviates from the correct answer and adopts an incorrect consensus, thus spreading the error further into the discussion. This is the most common mistake that can be made in a multi-agent discussion, even though they may have already had gotten the correct answer.
Qwen’s PRM report [46] improves the performance of the CoT by using the PRM to choose the result with the highest score (prm@n) in place of traditional majority voting (major@n). Furthermore, they proposed the Reward Guide Search (RGS) strategy. For each reasoning step of the COT, they sample multiple times and select the highest scoring step by PRM. This approach mitigates the challenge in answer selection, but severely loses the diversity of generated results, and for complex problems, it is very easy to fall into the wrong local optimal solution and fail to search for the correct answer.
@@ E.1 Prompt
For all the algorithms mentioned in the experiments, we used the same prompt, which follows the Qwen’s official recommendations. The prompt is displayed as follows:
@@
Obviously, we differed from the traditional MAS approach in that we ditched the role-play method altogether and didn’t use any fancy prompts. This enhances the robustness of the algorithm on different base models.
@@ E.2 More Details of Reason Step by Step
Therefore, to implement our algorithm, it is required that the base model has the ability of text completion or prefix continuation, and the ability to set stop words. Fortunately, all open-source models and most of the closed source models support the above abilities.
@@ Appendix G Further Discussion
In practice, however, evaluators often function as black boxes. Influenced by training data and specific architectures (e.g., Process Reward Models (PRMs) or Outcome Reward Models (ORMs)), their assessments offer imperfect guidance. Fitness scores might not accurately represent solution potential, and reliance on them risks premature convergence. Different evaluator types present distinct challenges: PRMs, focused on process correctness, might undervalue high-potential solutions with unconventional steps, while ORMs, assessing only final outcomes, lack process insight, potentially rewarding flawed reasoning or penalizing sound but imperfect processes. Both limitations hinder reasoning capability development. These inherent constraints (PRM oversight, ORM detail limits) can trap algorithms in local optima reflecting evaluator biases, diminishing diversity and impairing global exploration. Consequently, maintaining diversity while balancing exploitation/exploration is critical. Future algorithms must integrate diversity preservation mechanisms alongside fitness optimization to mitigate these challenges.
In summary, future research on agent-based swarm intelligence should prioritize two pivotal directions: enhancing efficient search for high-quality, near-optimal solutions and devising robust strategies for diversity maintenance to counteract black-box evaluator limitations (e.g., PRMs, ORMs) and reduce local optima susceptibility. Progress in these complementary areas is essential for advancing LLM performance in complex reasoning tasks.

### v2

@@ 2.1 Multi-Agent Systems
While the MAD scheme has shown effectiveness, previous systematic evaluations zhang2025if reveal that many MAD methods perform even worse than CoT methods. The advantage of solution diversity introduced by heterogeneous agent roles does not always hold, as the agents ultimately rely on the same underlying large language model, which leads to high similarities of the solutions generated by agents. Furthermore, the similarity between the solutions gradually increases during the collaboration process, since the LLM is progressively influenced by the accumulated context from historical interaction information. This factor makes it difficult to find the optimal results in the final despite the use of a collaborative scheme. Different from them, we propose integrating swarm intelligence algorithms into the reasoning process to enhance the discovery of optimal solutions. We conceptualize the LLM’s reasoning process as an exploration of the solution space and adopt heuristic algorithms to efficiently identify optimal solutions.
@@ 4 Methodology
Swarm intelligence algorithms typically explore the solution space via individual exploration and collective cooperation to identify the global optimum. This objective is analogous to that of LLM-based agents when undertaking reasoning tasks. Consequently, we propos a novel paradigm named ASI, which conceptualized the LLM’s reasoning process as an individual’s search for the global optimum within a solution space. Specifically, for a given problem $Q$, it owns a solution space $S$. Each dimension corresponds to a token, with its value derived from the LLM’s vocabulary. Every point within this solution space denotes a solution in the form of a reasoning path, and the essence of LLM reasoning is to generate such a path. Therefore, we apply swarm intelligence algorithms to LLM reasoning, which leverages the evaluator $E$ to evaluate the quality of the reasoning paths and guide the search, iteratively refining reasoning paths to efficiently and effectively uncover the global optima. Furthermore, based on the ASI paradigm, we propose the Swarm Intelligence Enhancing Reasoning (SIER) framework, which enhances solution space exploration by expanding the diversity of the search paths. This section first presents an overview of the SIER framework and detailed explanations of its key processes.
@@ B.1 Analysis of MAD Framework
Although the existing MAD framework is based on the hypothesis that LLMs can play different roles under various human prompts, unlike simple CoT, heterogeneous prompts can help agents collectively diverge their thinking and decide the final answer, which attempts to achieve two goals:
Brainstorming: Broaden the width of exploration of the solution space and obtain a variety of candidate solutions when facing complex problems.
@@
However, existing evaluation work suggests that MAD frameworks do not perform as well as expected in terms of the above two goals. [45] has experimentally found that many MAD frameworks perform worse than single-point methods and that self-consistency (i.e., by sampling the CoT results multiple times and voting for the best result) outperforms MAD frameworks significantly, as shown in Table 3.
@@
Besides, to analyze the reasons for poor MAD performance, [36] has identified two key problems that are relevant to the goal Selection: (1) Judgment Error: This occurs when the judger decides the final answer. If the answers vary between agents, the judger may choose the wrong option as the final verdict. This is especially true when the decision is made in a tie. (2) Wrong Answer Propagation: This error occurs when an agent, influenced by input from others, deviates from the correct answer and adopts an incorrect consensus, thus spreading the error further into the discussion. This is the most common mistake that can be made in a multi-agent discussion, even though they may have already gotten the correct answer.
Qwen’s PRM report [46] improves the performance of the CoT by using the PRM to choose the result with the highest score (prm@n) in place of traditional majority voting (major@n). Furthermore, they propose the Reward Guide Search (RGS) strategy. For each reasoning step of the COT, they sample multiple times and select the highest-score step by PRM. This approach mitigates the challenge in answer selection, but severely loses the diversity of generated results, and for complex problems, it is very easy to fall into the wrong local optimal solution and fail to search for the correct answer.
@@ E.1 Prompt
For all the algorithms mentioned in the experiments, we used the same prompt, which follows Qwen’s official recommendations. The prompt is displayed as follows:
@@
We differed from the traditional MAS approach in that we ditched the role-play method altogether and didn’t use any fancy prompts. This enhances the robustness of the algorithm on different base models.
@@ E.2 More Details of Reason Step by Step
Therefore, to implement our algorithm, it is required that the base model has the ability of text completion or prefix continuation, and the ability to set stop words. Fortunately, all open-source models and most of the closed-source models support the above abilities.
@@ F.1 Parameter Analysis
F.2 Tree Search Strategy - Reasoning via Planning
Reasoning via Planning (RAP) is a new reasoning framework for Large Language Models (LLMs) that aims to overcome the shortcomings of LLMs in generating task execution plans, and complex mathematical, logical, and common-sense reasoning.RAP does this by reorienting LLMs as world models and reasoning agents and combining them with planning algorithms based on Monte Carlo Tree Search for reasoning in a wide space of strategic exploration in a wide range of reasoning spaces. During the reasoning process, LLM as an agent gradually builds a reasoning tree guided by LLM (as a world model) and task-specific rewards, and obtains highly rewarding reasoning paths by striking an appropriate balance between exploration and exploitation. We refer to the experimental setup in this paper and obtain the experimental results shown in Fig. 6.
Figure 6: Performance comparison of SIER and RAP.
According to the experimental results, the SIER method proposed in this paper outperforms the RAP method on most datasets, especially on the GSM8K and LiveMathBench datasets. Specifically, the RAP method is inferior to SIER in terms of search efficiency and performance, probably due to the lack of reasonable sampling and pruning strategies. Moreover, RAP adjusts the values of the tree nodes through iterative optimization, and the model may concentrate on certain subtrees prematurely, i.e., it is easier to fall into local optima, which leads to difficulties in finding the global optima on complex problems. In contrast, the SIER method employs a more efficient search strategy and is able to explore the solution space better, resulting in higher performance.
@@ Appendix G Further Discussion
In practice, however, evaluators often function as black boxes. Influenced by training data and specific architectures (e.g., Process Reward Models (PRMs) or Outcome Reward Models (ORMs)), their assessments offer imperfect guidance. Fitness scores might not accurately represent solution potential, and reliance on them risks premature convergence. Different evaluator types present distinct challenges: PRMs, focused on process correctness, might undervalue high-potential solutions with unconventional steps, while ORMs, assessing only final outcomes, lack process insight, potentially rewarding flawed reasoning or penalizing sound but imperfect processes. Both limitations hinder reasoning capability development. These inherent constraints (PRM oversight, ORM detail limits) can trap algorithms in local optima reflecting evaluator biases, diminishing diversity, and impairing global exploration. Consequently, maintaining diversity while balancing exploitation/exploration is critical. Future algorithms must integrate diversity preservation mechanisms alongside fitness optimization to mitigate these challenges.
In summary, future research on agent-based swarm intelligence should prioritize two pivotal directions: enhancing the efficient search for high-quality, near-optimal solutions and devising robust strategies for diversity maintenance that counteract black-box evaluator limitations (e.g., PRMs, ORMs) and reduce local optima susceptibility. Progress in these complementary areas is essential for advancing LLM performance in complex reasoning tasks.